# Mitigating Hallucination of Large Vision-Language Models via Dynamic Logits Calibration

### v1

@@ 1 Introduction
The field of multimodal artificial intelligence has witnessed remarkable progress through the development of Large Vision-Language Models (LVLMs) [7, 12, 23, 26, 30, 42, 46, 56]. By integrating visual encoders [35, 51] with large language models (LLMs) [2, 10, 40, 41], contemporary LVLMs have successfully unified visual and textual information within a common semantic framework. These models demonstrate impressive capabilities across diverse visual tasks [3, 4, 21, 31], ranging from straightforward Visual Question Answering (VQA) to sophisticated scene reasoning.
@@
Despite their remarkable advances in multimodal comprehension, LVLMs remain plagued by a critical limitation: hallucination, the generation of text contradicting visual input [4, 5, 17, 22, 57]. This phenomenon introduces deceptive inconsistencies, as the model produces plausible but factually unfaithful descriptions of the observed scene. This issue is especially critical in high-stakes domains like medical diagnosis [6, 24] and autonomous systems [8, 34], where such inconsistencies may incur serious real-world consequence, further hindering the practical deployment of LVLMs.
@@
Figure 1: Comparison of decoding methods across three dimensions: Speed, Correctness, and Detailedness using LLaVA-1.5 [30] (max new tokens: 64), with all experiments conducted on a single NVIDIA A100 GPU. The vertical axis represents correctness, while the horizontal axis shows tokens per second (TPS). Bubble size corresponds to Detailedness of responses. Our method achieves superior accuracy and detailedness while maintaining decoding efficiency.
@@
To address the hallucination issue, numerous training-free decoding strategies have emerged. These approaches have garnered significant attention due to their independence from high-quality annotated datasets and resource-intensive computational requirements. Based on their intervention mechanisms, existing methods can be systematically categorized into three main paradigms: 1) Contrastive Decoding [18, 22, 43]; 2) Guided Decoding Frameworks [9, 13, 53]; and 3) Latent Space Intervention Methods [17, 32, 39, 57]. While these strategies demonstrate some effectiveness in generation tasks, systematic analysis reveals three fundamental limitations. First, static constraint mechanisms fail to dynamically adapt to semantic drift during generation, leaving the model vulnerable to hallucinations in later stages as linguistic priors dominate over visual grounding. Second, most methods require multiple forward passes or regeneration, drastically reducing inference efficiency. Third, rigid intervention rules lack contextual sensitivity, either enforcing overly strict visual alignment or allowing unchecked deviations, thereby distorting the balance between fidelity and creativity.
@@
Addressing these limitations-namely the lack of dynamic adaptation to context, the high inference cost, and the distortion of reality due to rigid rules-demands a deeper investigation. Our work identifies that a core issue underlying hallucination is the increasing reliance on linguistic priors during decoding, causing progressive detachment from visual evidence despite initial alignment. Motivated by this observation and the need to overcome the shortcomings of existing methods, we introduce Dynamic Logits Calibration (DLC) as an efficient and training-free decoding strategy. DLC employs the CLIP model as a dynamic probe to continuously monitor semantic correspondence between the image and generated text during token generation. It evaluates the semantic alignment of both the current text prefix and potential next tokens relative to the image, subsequently adjusting the logits to favor visually grounded options. This mechanism empowers tokens possessing a high "relative visual advantage" to supersede strong linguistic priors, thus diminishing the likelihood of visual hallucination while preserving generation coherence and maintaining high inference speed. As initially evidenced in Figure 1, DLC exhibits superior efficiency compared to other leading approaches. The concerned efficacy is validated through extensive experiments across multiple benchmarks, demonstrating its successful integration into diverse LVLM architectures such as LLaVA [30], InstructBLIP [12], and MiniGPT-4 [56]. In summary, our key contributions are threefold:
@@ 2 Related Work
Hallucination issue in LVLMs. The root cause of LVLM hallucination lies in the inherent tension between parametric knowledge dominance and visual evidence grounding: while users expect models to generate responses strictly conditioned on visual inputs, LVLMs frequently default to priors learned from their textual training corpus, favoring statistically likely but visually ungrounded outputs [15, 16, 22]. This over-reliance on linguistic patterns on top of pixel-accurate reasoning leads to hallucinations when the model’s generative biases override fidelity to the input image (see Appendix A.1 for additional background on LVLMs).
The prior approaches broadly fall into three paradigms. 1) Data-driven optimization [5, 19, 29, 45, 48, 49, 50, 54], including fine-grained contrastive learning, reinforcement learning, and curated instruction tuning with noisy or adversarial datasets. While striving for alignment enhancement, these methods demand expensively annotated data and often risk overfitting issue. 2) Post-hoc correction mechanisms [37, 44, 47, 52, 55], such as auxiliary reviser networks, image regeneration feedback, or external knowledge grounding. Though effective, they introduce significant inference latency and reliance on auxiliaries. 3) Decoding strategy refinement [9, 13, 17, 18, 22, 32, 39, 43, 53, 57], exemplified by VCD [22] that attenuates textual bias by contrasting original and distorted visual inputs, OPERA [17] that utilizes over-trust penalty and a retrospection allocation to alleviate aggregation patterns, and DoLa [11] that mitigates factual hallucinations via layer-wise logit calibration. Despite data dependency being avoided, these training-free methods exhibit critical limitations as discussed in the Introduction. DLC addresses these issues by employing real-time visual alignment assessment (Section 4.1) and adaptive logit modulation (Section 4.2), providing a more faithful decoding stage.
@@ 3.1 Generative Process of LVLMs
An LVLM, parameterized by $\theta$, processes jointly visual images $v$ and textual queries $x$. Typically, a pre-trained network encodes $v$ into visual features. An alignment module (e.g., Q-Former [26] or linear projection [30]) then maps these features into the textual semantic space. These aligned visual representations are concatenated with the embedded text query $x$ to form a joint input representation. Conditioned on this unified input, the LVLM auto-regressively predicts the next token $y_{t}$:
@@ 3.2 CLIP Model and CLIP Score
The Contrastive Language-Image Pre-training (CLIP) model [35], trained on extensive datasets of image-text pairs using a contrastive learning objective, learns to embed visual and textual information into a shared, high-dimensional semantic space. Formally, for a given image $v$ and text sequence $s$, we calculate their semantic similarity using cosine similarity, denoted as $\mathrm{CLIP}(v,s)$:
@@ 5.1 Experimental Settings
Evaluated LVLMs. We evaluate the effectiveness of our proposed DLC on three state-of-the-art LVLMs: LLaVA-1.5 [30], InstructBLIP [12] and MiniGPT-4 [56]. All benchmarked models adopt a dual-modality architecture, incorporating pre-trained vision encoders like CLIP [35] vision encoder and language models (e.g., Vicuna [10], LLaMA [41]) as fundamental components. Notably, all of models used in our paper are equipped with a 7B LLMs.
Benchmarks.To comprehensively evaluate our proposed method, we employed a suite of established benchmarks. These include:(1) CHAIR [36], for quantifying object hallucination in image captions;(2) GPT-4 Assisted Evaluations [54], to detect more fine-grained hallucination types;(3) GPT-4o assisted evaluations [47], for a comprehensive analysis of hallucinations and text quality; (4) LLaVA-Bench-in-the-Wild [30], on which we performed targeted case studies to demonstrate the efficacy of our proposed DLC in addressing complex tasks and adapting to novel domains. Detailed descriptions of each benchmark, are provided in Appendix A.4.
Baselines. To evaluate our training-free LVLM decoding strategies, we have compared them with standard decoding methods, including Nucleus Sampling, Greedy Sampling. We also assessed the performance of our method against several state-of-the-art decoding approaches: VCD [22], ICD [43], SID [18], and OPERA [17]. For comprehensive comparisons, we implemented VCD, ICD, SID, and DLC using both nucleus sampling (Top-p=1). The reported performance of these baselines is based on our re-implementation using their publicly available code.
@@ 5.2 Experimental Results
Max Token Methods LLaVA-1.5 InstuctBLIP MiniGPT-4
@@
CHAIR Evaluation. Following [17, 18], we randomly sample 500 images from COCO [28] validation and prompt the evaluated models with "Please describe this image in detail" to generate descriptions under two maximum new token limits. Moreover, to address CHAIR’s [36] image sensitivity, three distinct image sets were used for robust evaluation. Table 1 shows our method generally surpasses others, validating its efficacy. Notably, our method also exhibits superior generation speed over other elaborately designed approaches. Moreover, as our approach mitigates semantic drift, its superiority is particularly evident with a 512 Max Token setting.
@@
GPT-4 Assisted Evaluation. Following [17, 18, 54], we evaluate on 200 images from the VG dataset using the prompt "Please describe this image in detail," with a maximum of 512 new tokens generated. Additionally, we evaluated the model’s generation speed using tokens per second (TPS). As illustrated in Figure 5, we observe that our DLC achieves state-of-the-art results on metrics related to both hallucinated sentences and words. This indicates that DLC can effectively assist models in mitigating hallucination issues stemming from strong semantic priors. Consistent with observations for OPERA [17], we also find that for certain models, DLC may reduce the length of the generated text, which can be attributed to the reduction of superfluous hallucinatory content.
@@
GPT-4o Assisted Evaluation. To further analyze the hallucinations and text quality for open-end generation tasks, we employ GPT-4o as an advanced evaluation proxy. For evaluation, models generate detailed descriptions for 500 randomly selected COCO [28] images, prompted with: "Please describe this image in detail.". Comparative analyses are performed, contrasting our DLC with several established baselines, including Nucleus Sampling, VCD [22], OPERA [17], and SID [18]. The empirical results, summarized in Table 3, indicate that DLC consistently outperforms these baselines on most metrics. In particular, DLC demonstrates a significant improvement in correctness over conventional sampling-based decoding, while maintaining a comparable degree of detail. The close alignment of GPT-4o’s perceptual and reasoning faculties with human judgment lends substantial credibility to these findings, suggesting that the observed enhancements in hallucination reduction and text quality are indicative of genuine improvements from a human evaluative perspective (full GPT-4/GPT-4o prompt is provided in the Appendix A.9).
@@
Figure 5: GPT-4 assisted benchmark [54]. Six aspects of values are analyzed, including the number of sentences per image (SPI), the tokens per second (TPS), the number of hallucinated sentences per image (HSPI), the number of hallucinated words per image (HWPI), the ratio of hallucinated sentences (HSR), and the ratio of hallucinated words (HWR). Note that larger SPI and TPS, smaller HSPI, HWPI, HSR and HWR are better. Larger radar indicates better performance.
@@
Figure 6: Qualitative comparison on LLaVA-Bench-in-the-Wild [30]. Under identical sampling decoding strategy, our DLC approach is shown to significantly reduce hallucinatory content and enhance descriptive detail compared to baseline methods like Nucleus Sampling and VCD.
@@
Table 2: Assessing hallucination on COCO [28] with GPT-4o. Performance is assessed on correctness ($C$) and detailedness ($D$).
@@
Table 3: Assessing hallucination on LLaVA-Bench [30] with GPT-4o. Performance is assessed on correctness ($C$) and detailedness ($D$).
@@
Case Study on LLaVA-Bench. To complement our use of dedicated metrics like CHAIR [36] and evaluations facilitated by GPT-4/GPT-4o, we present a qualitative analysis of DLC’s hallucination mitigation capabilities through case studies from the LLaVA-Bench-in-the-Wild dataset [30]. As illustrated in Figure 6 (with additional cases in Appendix A.8), the DLC approach markedly reduces hallucinatory content and enhances descriptive detail under identical image and prompt conditions. These qualitative observations are further supported by a comprehensive GPT-4o evaluation of all LLaVA-Bench generation outputs, detailed in Table 3, which substantiates DLC’s superior performance over competing methods.

### v2

@@ 1 Introduction
The field of multimodal artificial intelligence has witnessed remarkable progress through the development of Large Vision-Language Models (LVLMs) (Chen et al., 2023; Dai et al., 2023; Li et al., 2024, 2023b; Liu et al., 2023; Wang et al., 2024a; Ye et al., 2024; Zhu et al., 2023). By integrating visual encoders (Radford et al., 2021; Zhai et al., 2023) with large language models (LLMs) (Bai et al., 2023a; Chiang et al., 2023; Touvron et al., 2023a, b), contemporary LVLMs have successfully unified visual and textual information within a common semantic framework. These models demonstrate impressive capabilities across diverse visual tasks (Bai et al., 2023b, 2024; Lai et al., 2024; Liu et al., 2024b), ranging from straightforward Visual Question Answering (VQA) to sophisticated scene reasoning.
@@
Despite their remarkable advances in multimodal comprehension, LVLMs remain plagued by a critical limitation: hallucination, the generation of text contradicting visual input (Bai et al., 2024; Chen et al., 2025; Huang et al., 2024; Leng et al., 2024; Zhuang et al., 2025). This phenomenon introduces deceptive inconsistencies, as the model produces plausible but factually unfaithful descriptions of the observed scene. This issue is especially critical in high-stakes domains like medical diagnosis (Chen et al., 2024a; Li et al., 2023a) and autonomous systems (Chen et al., 2024b; Nie et al., 2024), where such inconsistencies may incur serious real-world consequence, further hindering the practical deployment of LVLMs.
@@
Figure 1: Comparison of decoding methods across three dimensions: Speed, Correctness, and Detailedness using LLaVA-1.5 (Liu et al., 2023) (max new tokens: 64), with all experiments conducted on a single NVIDIA A100 GPU. The vertical axis represents correctness, while the horizontal axis shows tokens per second (TPS). Bubble size corresponds to Detailedness of responses. Our method achieves superior accuracy and detailedness while maintaining decoding efficiency.
@@
To address the hallucination issue, numerous training-free decoding strategies have emerged. These approaches have garnered significant attention due to their independence from high-quality annotated datasets and resource-intensive computational requirements. Based on their intervention mechanisms, existing methods can be systematically categorized into three main paradigms: 1) Contrastive Decoding (Huo et al., 2025; Leng et al., 2024; Wang et al., 2024b); 2) Guided Decoding Frameworks (Chen et al., 2024c; Deng et al., 2024; Zhao et al., 2024); and 3) Latent Space Intervention Methods (Huang et al., 2024; Liu et al., 2025; Tang et al., 2025; Zhuang et al., 2025). While these strategies demonstrate some effectiveness in generation tasks, systematic analysis reveals three fundamental limitations. First, static constraint mechanisms fail to dynamically adapt to semantic drift during generation, leaving the model vulnerable to hallucinations in later stages as linguistic priors dominate over visual grounding. Second, most methods require multiple forward passes or regeneration, drastically reducing inference efficiency. Third, rigid intervention rules lack contextual sensitivity, either enforcing overly strict visual alignment or allowing unchecked deviations, thereby distorting the balance between fidelity and creativity.
@@
Addressing these limitations-namely the lack of dynamic adaptation to context, the high inference cost, and the distortion of reality due to rigid rules-demands a deeper investigation. Our work identifies that a core issue underlying hallucination is the increasing reliance on linguistic priors during decoding, causing progressive detachment from visual evidence despite initial alignment. Motivated by this observation and the need to overcome the shortcomings of existing methods, we introduce Dynamic Logits Calibration (DLC) as an efficient and training-free decoding strategy. DLC employs the CLIP model as a dynamic probe to continuously monitor semantic correspondence between the image and generated text during token generation. It evaluates the semantic alignment of both the current text prefix and potential next tokens relative to the image, subsequently adjusting the logits to favor visually grounded options. This mechanism empowers tokens possessing a high "relative visual advantage" to supersede strong linguistic priors, thus diminishing the likelihood of visual hallucination while preserving generation coherence and maintaining high inference speed. As initially evidenced in Figure 1, DLC exhibits superior efficiency compared to other leading approaches. The concerned efficacy is validated through extensive experiments across multiple benchmarks, demonstrating its successful integration into diverse LVLM architectures such as LLaVA (Liu et al., 2023), InstructBLIP (Dai et al., 2023), and MiniGPT-4 (Zhu et al., 2023). In summary, our key contributions are threefold:
@@ 2 Related Work
Hallucination issue in LVLMs. The root cause of LVLM hallucination lies in the inherent tension between parametric knowledge dominance and visual evidence grounding: while users expect models to generate responses strictly conditioned on visual inputs, LVLMs frequently default to priors learned from their textual training corpus, favoring statistically likely but visually ungrounded outputs (Gupta et al., 2022; Han et al., 2022; Leng et al., 2024). This over-reliance on linguistic patterns on top of pixel-accurate reasoning leads to hallucinations when the model’s generative biases override fidelity to the input image (see Appendix A.1 for additional background on LVLMs).
The prior approaches broadly fall into three paradigms. 1) Data-driven optimization (Chen et al., 2025; Jiang et al., 2024; Liu et al., 2024a; Xiao et al., 2024; Yu et al., 2024a, b; Yue et al., 2024; Zhao et al., 2023), including fine-grained contrastive learning, reinforcement learning, and curated instruction tuning with noisy or adversarial datasets. While striving for alignment enhancement, these methods demand expensively annotated data and often risk overfitting issue. 2) Post-hoc correction mechanisms (Sahu et al., 2024; Wu et al., 2024; Yin et al., 2024; Zhang et al., 2025; Zhou et al., 2024), such as auxiliary reviser networks, image regeneration feedback, or external knowledge grounding. Though effective, they introduce significant inference latency and reliance on auxiliaries. 3) Decoding strategy refinement (Chen et al., 2024c; Deng et al., 2024; Huang et al., 2024; Huo et al., 2025; Leng et al., 2024; Liu et al., 2025; Tang et al., 2025; Wang et al., 2024b; Zhao et al., 2024; Zhuang et al., 2025), exemplified by VCD (Leng et al., 2024) that attenuates textual bias by contrasting original and distorted visual inputs, OPERA (Huang et al., 2024) that utilizes over-trust penalty and a retrospection allocation to alleviate aggregation patterns, and DoLa (Chuang et al., 2024) that mitigates factual hallucinations via layer-wise logit calibration. Despite data dependency being avoided, these training-free methods exhibit critical limitations as discussed in the Introduction. DLC addresses these issues by employing real-time visual alignment assessment (Section 4.1) and adaptive logit modulation (Section 4.2), providing a more faithful decoding stage.
@@ 3.1 Generative Process of LVLMs
An LVLM, parameterized by $\theta$, processes jointly visual images $v$ and textual queries $x$. Typically, a pre-trained network encodes $v$ into visual features. An alignment module (e.g., Q-Former (Li et al., 2023b) or linear projection (Liu et al., 2023)) then maps these features into the textual semantic space. These aligned visual representations are concatenated with the embedded text query $x$ to form a joint input representation. Conditioned on this unified input, the LVLM auto-regressively predicts the next token $y_{t}$:
@@ 3.2 CLIP Model and CLIP Score
The Contrastive Language-Image Pre-training (CLIP) model (Radford et al., 2021), trained on extensive datasets of image-text pairs using a contrastive learning objective, learns to embed visual and textual information into a shared, high-dimensional semantic space. Formally, for a given image $v$ and text sequence $s$, we calculate their semantic similarity using cosine similarity, denoted as $\mathrm{CLIP}(v,s)$:
@@ 5.1 Experimental Settings
Evaluated LVLMs. We evaluate the effectiveness of our proposed DLC on three state-of-the-art LVLMs: LLaVA-1.5 (Liu et al., 2023), InstructBLIP (Dai et al., 2023) and MiniGPT-4 (Zhu et al., 2023). All benchmarked models adopt a dual-modality architecture, incorporating pre-trained vision encoders like CLIP (Radford et al., 2021) vision encoder and language models (e.g., Vicuna (Chiang et al., 2023), LLaMA (Touvron et al., 2023b)) as fundamental components. Notably, all of models used in our paper are equipped with a 7B LLMs.
Benchmarks.To comprehensively evaluate our proposed method, we employed a suite of established benchmarks. These include:(1) CHAIR (Rohrbach et al., 2018), for quantifying object hallucination in image captions;(2) GPT-4 Assisted Evaluations (Zhao et al., 2023), to detect more fine-grained hallucination types;(3) GPT-4o assisted evaluations (Yin et al., 2024), for a comprehensive analysis of hallucinations and text quality; (4) LLaVA-Bench-in-the-Wild (Liu et al., 2023), on which we performed targeted case studies to demonstrate the efficacy of our proposed DLC in addressing complex tasks and adapting to novel domains. Detailed descriptions of each benchmark, are provided in Appendix A.4.
Baselines. To evaluate our training-free LVLM decoding strategies, we have compared them with standard decoding methods, including Nucleus Sampling, Greedy Sampling. We also assessed the performance of our method against several state-of-the-art decoding approaches: VCD (Leng et al., 2024), ICD (Wang et al., 2024b), SID (Huo et al., 2025), and OPERA (Huang et al., 2024). For comprehensive comparisons, we implemented VCD, ICD, SID, and DLC using both nucleus sampling (Top-p=1). The reported performance of these baselines is based on our re-implementation using their publicly available code.
@@ 5.2 Experimental Results
Max
Token
Methods LLaVA-1.5 InstuctBLIP MiniGPT-4
@@
CHAIR Evaluation. Following (Huang et al., 2024; Huo et al., 2025), we randomly sample 500 images from COCO (Lin et al., 2014) validation and prompt the evaluated models with "Please describe this image in detail" to generate descriptions under two maximum new token limits. Moreover, to address CHAIR’s (Rohrbach et al., 2018) image sensitivity, three distinct image sets were used for robust evaluation. Table 1 shows our method generally surpasses others, validating its efficacy. Notably, our method also exhibits superior generation speed over other elaborately designed approaches. Moreover, as our approach mitigates semantic drift, its superiority is particularly evident with a 512 Max Token setting.
@@
GPT-4 Assisted Evaluation. Following (Huang et al., 2024; Huo et al., 2025; Zhao et al., 2023), we evaluate on 200 images from the VG dataset using the prompt "Please describe this image in detail," with a maximum of 512 new tokens generated. Additionally, we evaluated the model’s generation speed using tokens per second (TPS). As illustrated in Figure 5, we observe that our DLC achieves state-of-the-art results on metrics related to both hallucinated sentences and words. This indicates that DLC can effectively assist models in mitigating hallucination issues stemming from strong semantic priors. Consistent with observations for OPERA (Huang et al., 2024), we also find that for certain models, DLC may reduce the length of the generated text, which can be attributed to the reduction of superfluous hallucinatory content.
@@
GPT-4o Assisted Evaluation. To further analyze the hallucinations and text quality for open-end generation tasks, we employ GPT-4o as an advanced evaluation proxy. For evaluation, models generate detailed descriptions for 500 randomly selected COCO (Lin et al., 2014) images, prompted with: "Please describe this image in detail.". Comparative analyses are performed, contrasting our DLC with several established baselines, including Nucleus Sampling, VCD (Leng et al., 2024), OPERA (Huang et al., 2024), and SID (Huo et al., 2025). The empirical results, summarized in Table 3, indicate that DLC consistently outperforms these baselines on most metrics. In particular, DLC demonstrates a significant improvement in correctness over conventional sampling-based decoding, while maintaining a comparable degree of detail. The close alignment of GPT-4o’s perceptual and reasoning faculties with human judgment lends substantial credibility to these findings, suggesting that the observed enhancements in hallucination reduction and text quality are indicative of genuine improvements from a human evaluative perspective (full GPT-4/GPT-4o prompt is provided in the Appendix A.12).
@@
Figure 5: GPT-4 assisted benchmark (Zhao et al., 2023). Six aspects of values are analyzed, including the number of sentences per image (SPI), the tokens per second (TPS), the number of hallucinated sentences per image (HSPI), the number of hallucinated words per image (HWPI), the ratio of hallucinated sentences (HSR), and the ratio of hallucinated words (HWR). Note that larger SPI and TPS, smaller HSPI, HWPI, HSR and HWR are better. Larger radar indicates better performance.
@@
Figure 6: Qualitative comparison on LLaVA-Bench-in-the-Wild (Liu et al., 2023). Under identical sampling decoding strategy, our DLC approach is shown to significantly reduce hallucinatory content and enhance descriptive detail compared to baseline methods like Nucleus Sampling and VCD.
@@
Table 2: Assessing hallucination on COCO (Lin et al., 2014) with GPT-4o. Performance is assessed on correctness ($C$) and detailedness ($D$).
@@
Table 3: Assessing hallucination on LLaVA-Bench (Liu et al., 2023) with GPT-4o. Performance is assessed on correctness ($C$) and detailedness ($D$).
@@
Case Study on LLaVA-Bench. To complement our use of dedicated metrics like CHAIR (Rohrbach et al., 2018) and evaluations facilitated by GPT-4/GPT-4o, we present a qualitative analysis of DLC’s hallucination mitigation capabilities through case studies from the LLaVA-Bench-in-the-Wild dataset (Liu et al., 2023). As illustrated in Figure 6 (with additional cases in Appendix A.13), the DLC approach markedly reduces hallucinatory content and enhances descriptive detail under identical image and prompt conditions. These qualitative observations are further supported by a comprehensive GPT-4o evaluation of all LLaVA-Bench generation outputs, detailed in Table 3, which substantiates DLC’s superior performance over competing methods.